OpenStudy (stormswan):

Please help me! I will give a medal! Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (johnweldon1993):

Hint: Parallel lines have _____ slopes ...?

OpenStudy (stormswan):

well, identical....

OpenStudy (johnweldon1993):

Right, So we know the line we will have will also have a slope of -1 So automatically we are down to either B or C

OpenStudy (stormswan):

-1? doesn't it say -2?

OpenStudy (johnweldon1993):

Not quite...the line we are given is \[\large y = -x - 4\] if we compare that to slope intercept form... \[\large y = mx + b\] 'm' is our slope there...which is -1

OpenStudy (johnweldon1993):

Does that make sense?

OpenStudy (johnweldon1993):

Now, the line must pass through the point (-3,1) ...since we know the slope is to be -1...we can plug that slope and that point into the formula \(\large y = mx + b\) to solve for 'b' (our y-intercept) So we have \[\large 1 = (-1)(-3) + b\] and solve that for 'b'

OpenStudy (stormswan):

I don't understand how i would solve for b though.

OpenStudy (stormswan):

because wouldn't you have to do everything on both sides..?

OpenStudy (johnweldon1993):

Well lets see.. \[\large 1 = (-1)(-3) + b\] What is -1 times -3? it is 3 ...so we have \[\large 1 = 3 + b\] Now we want to solve for 'b'...so subtract 3 from both sides of the equation \[\large b = -2\]

OpenStudy (stormswan):

ooooh! ok that makes a lot more sense now, because i didn't know that i had to multiply those two negatives first.

OpenStudy (johnweldon1993):

What we now have...is the y-intercept of the line we wanted...so now...we can write what the equation of the parallel line is... \[\large y = -x - 2\]

OpenStudy (stormswan):

thank you so much! it totally makes sense now, you are awesome!

OpenStudy (johnweldon1993):

No problem :)